OpenStudy (anonymous):

What is the solution of |-1+x|=5

OpenStudy (anonymous):

solve for -1+x = 5 and then also for -1+x = -5

OpenStudy (anonymous):

-1+x=5 or -1+x=-5 -1+x=5 x=6 -1+x=-5 x=-4

OpenStudy (anonymous):

|-1+x|=5 solve for x x=-4 because this is an absolute value question, only positive numbers are included in our answer, there-fore: x=4

OpenStudy (anonymous):

|-1+x|=5 x=-4 or x=6

OpenStudy (anonymous):

sorry not -4 but 4